Grade 8General Physics

A ball thrown vertically upward with a velocity of 40 m/s.Calculate maximum height and time to return to the thrower.

Kundan chaudhary
8 Years agoGrade 8

1 Answer

Khimraj
8 Years ago
to calculate maximum height use third equation of motion
v2 = u2 +2as
0 = 402 + 2*(-10)*h
h = 80m
time to reach max height
v = u – gt
0 = 40 -10t
t = 4 sec
time to return to the thrower
t = 4*2 = 8 sec.
Hope it clears. If you like answer please approve it.